# V2X-DGPE: Addressing Domain Gaps and Pose Errors for Robust Collaborative 3D Object Detection

### v1

@@ 4.5 Pose Errors Reslut
5 Conclusion
In this paper, we propose a novel collaborative perception framework, V2X-DGPE, designed to address the domain gaps between heterogeneous nodes and pose errors. The framework leverages a Knowledge Distillation Framework and a Feature Compensation Module to learn domain-invariant representations from multi-source data. By integrating historical information as supplementary input, the model enables a more comprehensive understanding of the current scene. The Collaborative Fusion Module employs a heterogeneous self-attention mechanism to extract and integrate heterogeneous representations from vehicle and infrastructure, optimizing the transmission and aggregation. Additionally, a deformable attention mechanism is introduced to address pose errors, allowing the model to adaptively focus on critical parts of the input features by dynamically offsetting sampling points. Extensive experiments on the real-world DAIR-V2X dataset demonstrate the effectiveness of V2X-DGPE in addressing domain gaps and pose errors.
6 Acknowledge
This work was supported in part by the National Natural Science Foundation of China under Grant.

### v2

@@ IV-E Pose Errors Reslut
V CONCLUSION
In this paper, we propose a novel framework, V2X-DGPE, for robust collaborative perception. V2X-DGPE effectively reduces domain gaps between heterogeneous nodes and achieves high accuracy. V2X-DGPE outperforms SOTA intermediate fusion methods DI-V2X[12] by 1.1%/3.3% for AP@0.5/0.7. Furthermore, V2X-DGPE demonstrates strong robustness in handling pose error noises. Under various pose noise levels of Gaussian and Laplace noise, V2X-DGPE achieves state-of-the-art performance. In future works, we will leverage more historical data and extend the integration of multimodal sensors to enhance V2X collaborative perception and prediction.